vhong is 23 and denice is 15. how many years ago was vhong three times as old as aenice?
Answer by robertb(5830)

Let x = # of years ago...
==>23 - x = vhong's age x years ago and 15 - x = denice's age x years ago.
==> 23 - x = 3(15 - x)
23 - x = 45 - 3x
2x = 22 ==> x = 11.


P.S. This problem made me hungry. Gonna get some "foods"...